# Complexity of deep computations via topology of function spaces

### v1

@@ A Compositional Computation Structure (CCS) is a triple $(L,\mathcal{P},\Gamma)$, where
The elements of the semigroup $\Gamma$ are called the computations of the structure $(L,\mathcal{P},\Gamma)$.
If $\Delta\subseteq\Gamma$, we say that $\Delta\subseteq\Gamma$ is $R$-confined if $\gamma|_{L[r_{\bullet}]}:L[r_{\bullet}]\to L[r_{\bullet}]$ for every $r_{\bullet}\in R$ and $\gamma\in\Delta$. Elements in $\overline{\Gamma}\subseteq\mathcal{L}_{\text{sh}}$ are called (real-valued) deep computations or ultracomputations.
For a CCS $(L,\mathcal{P},\Gamma)$, we regard the elements of $\Gamma$ as “standard” finitary computations and the elements of $\overline{\Gamma}$, i.e., deep computations, as possibly infinitary limits of standard computations. The main goal of this paper is to study the computability, definability and computational complexity of deep computations. Since ultracomputations are defined through a combination of topological concepts (namely, topological closure) and structural and model-theoretic concepts (namely, models and types), we will import technology from both topology and model theory.
@@ 2.2. Computability and definability of deep computations and the Extendibility Axiom
Let $f:L\to\mathcal{L}$ be a function that maps each input state type $(P(v))_{P\in\mathcal{P}}\in\mathbb{R}^{\mathcal{P}}$ into an output state type $(P\circ f(v))_{P\in\mathcal{P}}\in\mathbb{R}^{\mathcal{P}}$.
@@
    We will say that $f$ is definable if for each $Q\in\mathcal{P}$, the output feature $Q{\circ}f:L\to\mathbb{R}$ is a definable predicate in the following sense: There is an approximating function $\varphi_{Q,\,K,\varepsilon}:L\to\mathcal{L}$ that can be built recursively out of a finite number of the (primitively computable) predicates in $\mathcal{P}$ by a finite number of applications of the finitary lattice operations $\land$ (=min) and $\lor$ (=max), the operations of $\mathbb{R}^{\mathbb{R}}$ as a vector algebra (that is, vector addition and multiplication and scalar multiplication) and the operators sup and inf applied on individual variables from $L$, and such that
@@
    Remark: What we have defined above is a model-theoretic concept; it is a special case of the concept of first-order definability for real-valued predicates in the model the theory of real-valued structures first introduced in [Iovino:1994] for model theory of functional analysis and now standard in model theory (see [Keisler:2023]). The $\land$ (=min) and $\lor$ (=max) operations correspond to the positive Boolean logical connectives “and” and “or”, and the sup and inf operators correspond to the first-order quantifiers, $\forall$ and $\exists$.
@@
It is shown in [alva2024approximability] that:
@@
We say that a CCS $(L,\mathcal{P},\Gamma)$ satisfies the Extendability Axiom if for all $\gamma\in\Gamma$, there is $\tilde{\gamma}:\mathcal{L}_{\text{sh}}\to\mathcal{L}_{\text{sh}}$ such that for every sizer $r_{\bullet}$ there is a sizer $s_{\bullet}$ such that $\tilde{\gamma}|_{\mathcal{L}[r_{\bullet}]}:\mathcal{L}[r_{\bullet}]\to\mathcal{L}[s_{\bullet}]$ is continuous. We refer to $\tilde{\gamma}$ as a free extension of $\gamma$.
@@
By the preceding remarks, the Extendability Axiom says that the elements of the semigroup $\Gamma$ are finitary computations. For the rest of the paper, fix for each $\gamma\in\Gamma$ a free extension $\tilde{\gamma}$ of $\gamma$. For any $\Delta\subseteq\Gamma$, let $\tilde{\Delta}$ denote $\{\tilde{\gamma}\colon\gamma\in\Delta\}$.
For a more detailed discussion of the Extendability Axiom, we refer the reader to [alva2024approximability].
For an illustrative example, we can frame Newton’s polynomial root approximation method in the context of a CCS (see Example 5.6 of [alva2024approximability] for details) as follows. Begin by considering the extended complex numbers $\hat{\mathbb{C}}$$\mathbb{C}\cup\{\infty\}$ with the usual Riemann sphere topology that makes it into a compact space (where unbounded sequences converge to $\infty$). In fact, not only is this space compact, but it is covered by the shard given by the sizer $(1,1,1)$ (the unit sphere is contained in the cube $1,1]^{3}$). The space $\hat{\mathbb{C}}$ is homeomorphic to the usual unit sphere $S^{2}:=\{(x,y,z):x^{2}+y^{2}+z^{2}=1\}$ of $\mathbb{R}^{3}$, by means of the stereographic projection and its inverse $\hat{\mathbb{C}}\to S^{2}$. This function is regarded as a triple of predicates $x,y,z:\hat{\mathbb{C}}\to[-1,1]$ where each will map an extended complex number to its corresponding real coordinate on the cube $1,1]^{3}$. Now fix the cubic complex polynomial $p(s):=s^{3}-1$, and consider the map which performs one step in Newton’s method at a particular (extended) complex number $s$, for finding a root of $p$, $\gamma_{p}:\hat{\mathbb{C}}\to\hat{\mathbb{C}}$. The explicit inner workings of $\gamma_{p}$ are irrelevant for this example, except for the fact that it is a continuous mapping. It follows that $(S^{2},\{x,y,z\},\{\gamma_{p}^{k}:k\in\mathbb{N}\})$ is a CCS. The idea is that repeated applications of $\gamma_{p}(s),\gamma_{p}\circ\gamma_{p}(s),\gamma_{p}\circ\gamma_{p}\circ\gamma_{p}(s),\dots$ would approximate a root of $p$ provided $s$ was a good enough initial guess.

### v2

@@ A Compositional Computation Structure (CCS) is a triple $(L,\mathcal{P},\Gamma)$, where
Elements of the semigroup $\Gamma$ are called the computations of the structure $(L,\mathcal{P},\Gamma)$. We assume that the identity map id on $L$ is an element of $\Gamma$ (which is thus not merely a semigroup but a monoid of transformations of $L$).
We topologize $\Gamma$ as a subset of the topological product $L^{L}$, where the “exponent” $L$ serves merely as an index set, but the “base” $L$ is topologized by type; consequently, one may identify $\Gamma$ with a subset of the topological product $(\mathbb{R}^{\mathcal{P}})^{L}$. More specifically, $\Gamma$ is identified with a subset of $\mathcal{L}^{L}$, which is a closed subspace of $(\mathbb{R}^{\mathcal{P}})^{L}$. Therefore, we have an inclusion $\overline{\Gamma}\subseteq\mathcal{L}^{L}$. Elements $\xi\in\overline{\Gamma}$ are called (real-valued) deep computations or ultracomputations.
A collection $R$ of sizers is exhaustive if $L=\bigcup_{r_{\bullet}\in R}L[r_{\bullet}]$ (shards $L[r_{\bullet}]$ exhaust $L$). A transformation $\gamma\in\Gamma$ is $R$-confined if $\gamma$ restricts to a map $\gamma|_{L[r_{\bullet}]}:L[r_{\bullet}]\to L[r_{\bullet}]$ (into $L[r_{\bullet}]$ itself) for every $r_{\bullet}\in R$. A subset $\Delta\subseteq\Gamma$ is $R$-confined if each $\gamma\in\Delta$ is.
Proposition 2.4.
If $\Delta\subseteq\Gamma$ is confined by an exhaustive sizer collection, then $\overline{\Delta}$ is a compact subset of $\mathcal{L}_{\mathrm{sh}}^{L}$.
Proof.
Assume that $R$ confines $\Delta$. For each $v\in L$, let $r_{\bullet}^{(v)}\in R$ be a sizer such that $v\in L[r_{\bullet}^{(v)}]$. An arbitrary $\gamma\in\Delta$ restricts to a map $\gamma\restriction L[r_{\bullet}^{(v)}]:L[r_{\bullet}^{(v)}]\to L[r_{\bullet}^{(v)}]$, so $\Gamma\subseteq K\coloneqq\prod_{v\in L}\mathcal{L}[r_{\bullet}^{(v)}]$. The space $K$ is a product of compact spaces, hence compact, so $\overline{\Gamma}$ is a closed, hence compact subset thereof, and a subset of $\mathcal{L}_{\mathrm{sh}}^{L}\supseteq K$ a fortiori. ∎
For a CCS $(L,\mathcal{P},\Gamma)$, we regard the elements of $\Gamma$ as “standard” finitary computations, and the elements of $\overline{\Gamma}$, i.e., deep computations, as possibly infinitary limits of standard computations. The main goal of this paper is to study the computability, definability and computational complexity of deep computations. Since ultracomputations are defined through a combination of topological concepts (namely, topological closure) and structural and model-theoretic concepts (namely, models and types), we will import technology from both topology and model theory.
@@ 2.2. Computability and definability of deep computations and the Extendibility Axiom
Let $f:L\to\mathcal{L}$ be a function that maps each input state type $(P(v))_{P\in\mathcal{P}}\in\mathbb{R}^{\mathcal{P}}$ to an output state type $(P\circ f(v))_{P\in\mathcal{P}}\in\mathbb{R}^{\mathcal{P}}$.
@@
    We will say that $f$ is definable if for each $Q\in\mathcal{P}$, the output feature $Q{\circ}f:L\to\mathbb{R}$ is a definable predicate in the following sense: There is an approximating function $\varphi_{Q,\,K,\varepsilon}:L\to\mathcal{L}$ that can be built recursively out of a finite number of the (primitively computable) predicates in $\mathcal{P}$ and by a finite number of iterations of the finitary lattice operations $\land$ (=min) and $\lor$ (=max), the operations of $\mathbb{R}^{\mathbb{R}}$ as a vector algebra (that is, vector addition and multiplication and scalar multiplication) and the operators sup and inf applied on individual variables from $L$, and such that
@@
    Remark: What we have defined above is a model-theoretic concept; it is a special case of the concept of first-order definability for real-valued predicates in the model the theory of real-valued structures first introduced in [IOV94] for model theory of functional analysis and now standard in model theory (see [KEI03]). The $\land$ (=min) and $\lor$ (=max) operations correspond to the positive Boolean logical connectives “and” and “or”, and the sup and inf operators correspond to the first-order quantifiers, $\forall$ and $\exists$.
@@
It is shown in [ADI+24] that:
@@
We say that a CCS $(L,\mathcal{P},\Gamma)$ satisfies the Extendibility Axiom if for all $\gamma\in\Gamma$, there is $\tilde{\gamma}:\mathcal{L}_{\text{sh}}\to\mathcal{L}_{\text{sh}}$ such that for every sizer $r_{\bullet}$ there is a sizer $s_{\bullet}$ such that $\tilde{\gamma}|_{\mathcal{L}[r_{\bullet}]}:\mathcal{L}[r_{\bullet}]\to\mathcal{L}[s_{\bullet}]$ is continuous. We refer to $\tilde{\gamma}$ as a free extension of $\gamma$.
@@
By the preceding remarks, the Extendibility Axiom says that the elements of the semigroup $\Gamma$ are finitary computations. For the rest of the paper, fix for each $\gamma\in\Gamma$ a free extension $\tilde{\gamma}$ of $\gamma$. For any $\Delta\subseteq\Gamma$, let $\tilde{\Delta}$ denote $\{\tilde{\gamma}\colon\gamma\in\Delta\}$.
For a more detailed discussion of the Extendibility Axiom, we refer the reader to [ADI+24].
2.3. Newton’s method as a CCS
Let $p(z)$ be a non-constant polynomial with complex coefficients. Newton’s method is an iterative method that is used to approximate a root of $p(z)$. Define the Newton map as:
$N_{p}(z)=z-\frac{p(z)}{p^{\prime}(z)}$
The method consists of taking an initial guess $z_{0}\in\mathbb{C}$ and iterating the rational map $N_{p}$ to obtain a sequence given by
$z_{n+1}=N_{p}(z_{n})=z_{n}-\frac{p(z_{n})}{p^{\prime}(z_{n})}$
For each root $r\in\mathbb{C}$ of $p(z)$, there exists $\varepsilon>0$ such that for any initial guess $z_{0}$ in the $\varepsilon$-ball centered at $r$, Newton’s iteration converges to $r$ (provided $p^{\prime}(r)\neq 0$) and the convergence is quadratic in that case, meaning the error at each step is roughly squared, causing the number of correct digits to double, leading to fast convergence.
Denote by $N_{p}^{n}:=N_{p}\circ N_{p}\circ\cdots\circ N_{p}$ (nth iteration of $N_{p}$). Given a root $r$ of $p(z)$, the set $B_{r}=\{z\in\mathbb{C}:\lim_{n\rightarrow\infty}N_{p}^{n}(z)=r\}$ is an open set called the basin of $r$. However, Newton’s method can fail to converge to any root for some choices of $z_{0}$. For example, consider the polynomial $p(z)=z^{3}-2z+2$. The Newton map is given by
$N_{p}(z)=z-\frac{z^{3}-2z+2}{3z^{2}-2}=\frac{2z^{3}-2}{3z^{2}-2}$
Notice that taking $z_{0}=0$ as an initial guess will yield the sequence $0,1,0,1,0,1,\dots$ that oscillates between 0 and $1$ but none of them are roots of $p(z)$. Another more chaotic way Newton’s method can fail to converge is when the sequence of iterations has no convergent subsequence. The set of such points is known as the Julia set associated to $N_{p}$ and it is typically a fractal. This can be visualizaed by adding a dash of color: let us give each complex number $z_{0}$ a color $(R,G,B)$ where $R,G,B\in[0,1]$ (so that $(1,0,0)$ is red, $(0,1,0)$ is green, $(0,0,1)$ is blue and $(0.5,0,0.5)$ is a light purple, for example). The values of $R$, $G$ and $B$ are determined by looking at the image of said number at each stage of the iteration, $N_{p}^{n}(z_{0})$, and computing the current distance to each of the roots of $p(z)$; so $R=1/d_{r}$ where $d_{r}$ is the positive distance to the root which is colored red, and so on. In this way, the roots themselves are colored red, green, and blue, and every other point gets a mix of the three colors. As the number of iterations increases, each point gets a sharper color, as the sequence of images $\{N_{p}^{n}(z_{0})\}_{n=1}^{\infty}$ converges to one of the three roots. Each stage, the complex plane looks as if out of focus because the coloring function is continuous. As the reader can see in Figure 1, the points at the boundary of each color class form the famous Newton fractal (of which, interestingly, Newton was unaware).
(a) After 1 iteration
(b) After 3 iterations
(c) After 100 iterations
Figure 1. Newton’s method approximating $p(z)=z^{3}-2z+2$. Notice the regions of divergence.
Another example of a Newton’s fractal is for $p(z)=z^{3}-1$. The roots of $p(z)$ are the 3rd roots of unity and the Newton map is given by:
$N_{p}(z)=z-\frac{z^{3}-1}{3z^{2}}=\frac{2z^{3}+1}{3z^{2}}$
In this case, there are three basins of attraction (one for each root) and the complement of their union is the Julia set, i.e., the common boundary.
(a) After 1 iteration
(b) After 3 iterations
(c) After 100 iterations
Figure 2. Newton’s method approximating $p(z)=z^{3}-1$.
We can frame Newton’s method as a CCS (satisfying the Extendibility Axiom) as follows. Given a non-constant polynomial $p(z)$, let $L\subseteq\mathbb{C}$ be the complement of the Julia set (this is called the Fatou set) associated to $N_{p}$. This is the set where Newton’s method is tame. By [BLA84, Corollary 4.6], $L$ is open and dense in $\mathbb{C}$. Let $\mathcal{P}=\{P_{1},P_{2},P_{3}\}$ where the map $z\mapsto(P_{1}(z),P_{2}(z),P_{3}(z))$ is the stereographic projection into the Riemann Sphere $S^{2}$, i.e.,
$P_{1}(z)=\frac{2{\rm Re}(z)}{|z|^{2}+1},$
$P_{2}(z)=\frac{2{\rm Im}(z)}{|z|^{2}+1},$
$P_{3}(z)=\frac{|z|^{2}-1}{|z|^{2}+1}.$
Since $L$ is dense in the extended complex plane, then $\mathcal{L}=S^{2}$. Computations are the iterations of Newton’s method, i.e., $\Delta=\Gamma=\{N_{p}^{n}:n\in\mathbb{N}\}$. Then, $(L,\mathcal{P},\Gamma)$ is a CCS. Since all computations $N_{p}^{n}$ are rational maps, they can be continuously extended to the extended complex plane $\overline{\mathbb{C}}=\mathbb{C}\cup\{\infty\}$, i.e., to $\mathcal{L}$. In particular, $(L,\mathcal{P},\Gamma)$ satisfy the Extendibility Axiom. The set of deep computations $\overline{\Delta}$ might behave different for various polynomials. Let us look at various examples:
Example 2.6.
[…]
$P_{1}(x)=\frac{2x}{x^{2}+1},$
$P_{2}(x)=\frac{x^{2}-1}{x^{2}+1}.$
The Newton’s method map $N_{p}:L\rightarrow L$ is given by
$N_{p}(x)=\frac{x^{2}+a}{2x}.$
Note that $\mathcal{L}=S^{1}$ and that each iterate $N_{p}^{n}$ can be continuously extended to the extended real line $\mathbb{R}\cup\{\infty\}$, i.e., $\mathcal{L}$. For example,
$\tilde{N_{p}}(x)=\begin{cases}\frac{x^{2}+a}{2x},&\text{if }x\in L;\\ \infty,&\text{if }x=0,\infty.\end{cases}$
For every initial guess $x\in L$, the limit $f(x)=\lim_{n\rightarrow\infty}N_{p}^{n}(x)$ converges pointwise to one of the roots. Moreover,
$f(x)=\begin{cases}\sqrt{a},&\text{if }x>0;\\ -\sqrt{a},&\text{if }x<0.\end{cases}$
Notice that $f$ can be extended to $\mathcal{L}$ by
$\tilde{f}(x)=\begin{cases}\sqrt{a},&\text{if }x>0;\\ -\sqrt{a},&\text{if }x<0;\\ \infty,&\text{if }x=0,\infty.\end{cases}$
However, $\tilde{f}$$\mathcal{L}\rightarrow\mathcal{L}$ is not continuous. The set $\overline{\tilde{\Delta}}$ of deep computations is $\tilde{\Delta}\cup\{\tilde{f}\}\subseteq B_{1}(\mathcal{L},\mathcal{L})$.
Example 2.7.
Newton’s method for $p(z)=z^{3}-2z+2$. Let $r_{1}$, $r_{2}$ and $r_{3}$ be the three roots of $p(z)$. Let $B_{1}$, $B_{2}$ and $B_{3}$ be their respective basins. Let $B$ be the basin of the attractive cycle $0,1,0,1,\dots$. Let $L=B\cup\bigcup_{i=1}^{3}B_{i}$ and $J$ be its complement, i.e., the Julia set. Notice that $N_{p}^{n}$ does not converge pointwise. However, the subsequences $N_{p}^{2n}$ and $N_{p}^{2n+1}$ are pointwise convergent to functions $f_{1}$ and $f_{2}$ respectively. $f_{1}$ and $f_{2}$ are two distinct deep computations. Note that for $z\in J$, no subsequence of $\tilde{N_{p}^{n}}(z)$ converges to a complex number. However, since $\mathcal{L}=S^{2}$ is compact there is a subsequence of $\tilde{N_{p}^{n}}(z)$ that converges to $\infty$. We can extend $f_{i}:L\rightarrow\mathcal{L}$ to $\tilde{f_{i}}$$\mathcal{L}\rightarrow\mathcal{L}$ by:
$\tilde{f_{i}}(z)=\begin{cases}f_{i}(z),&\text{if }z\in L;\\ \infty,&\text{if }z\in J.\end{cases}.$
Again, note that $\tilde{f_{i}}$ for $i=1,2$ are not continuous and that $\tilde{f_{i}}\in\overline{\tilde{\Delta}}$.
2.4. Finite precision threshold classifiers as a CCS
Let $L=2^{\mathbb{N}}$, i.e., the set consisting of all infinite binary sequences with the topology of pointwise convergence. Let $\mathcal{P}=\{P_{n}:n\in\mathbb{N}\}$ is the collection of projections, i.e., $P_{n}(x)=x(n)$ for each $n$ and each $x\in L$. In other words, $P_{n}$ is the predicate that reads the $n$th digit of a sequence. Notice that $L\subseteq\mathbb{R}^{\mathcal{P}}$ is closed. Therefore, $\mathcal{L}=L$. We denote by $0^{\infty}$ the infinite binary sequence consisting of 0s, and by $1^{\infty}$ the infinite binary sequence consisting of $1$s. The set of finite binary strings is denoted by $2^{<\mathbb{N}}$. Given a finite binary string $w$, we consider the transition $\phi_{w}:L\to L$ given by the rule
$\phi_{w}(x)=\begin{cases}1^{\infty},&\text{if }x|_{|w|}\leq_{\mathrm{lex}}w;\\ 0^{\infty},&\text{otherwise};\end{cases}$
where $|w|$ is the length of the string $w$ and $x|_{|w|}$ is the prefix of $x$ of length $|w|$. That is, $\phi_{w}(x)$ is equal to the constant sequence of ones if $x|_{|w|}$ comes before or is equal to $w$ in the lexicographic order of strings, and it is equal to the constant sequence of zeros otherwise. In words, $\phi_{w}$ checks if a number is less than or equal to the scalar value of threshold $w$ (the string $w$ is finite, hence the classifier has finite precision). Note that $P_{n}\circ\phi_{w}(x)=1$ if and only if $x|_{|w|}$ comes before $w$.
Proposition 2.8.
$\phi_{w}:2^{\mathbb{N}}\rightarrow 2^{\mathbb{N}}$ is continuous for all $w\in 2^{<\mathbb{N}}$.
Proof.
It suffices to prove that $P_{n}\circ\phi_{w}:2^{\mathbb{N}}\rightarrow\{0,1\}$ is continuous for all $n\in\mathbb{N}$. For simplicity, let us call $f:=P_{n}\circ\phi_{w}$, i.e.,
$f(x)=\begin{cases}1,&\text{if }x|_{|w|}\leq_{\mathrm{lex}}w;\\ 0,&\text{otherwise}.\end{cases}$
We first prove that $f^{-1}(1)=\{x\in 2^{\mathbb{N}}:x|_{|w|}\leq_{\mathrm{lex}}w\}$ is an open set. Fix $x_{0}\in f^{-1}(1)$. Let $t:=x_{0}|_{|w|}$ and consider the basic open set $t]=\{x\in 2^{\mathbb{N}}:x|_{|t|}=t\}$. Then, it is not difficult to check that $x_{0}\in[t]\subseteq f^{-1}(1)$. The same reasoning shows that $f^{-1}(0)$ is open. ∎
Let $\Delta=\{\phi_{w}:w\in 2^{<\mathbb{N}}\}\cup\{\mathbf{0^{\infty}},\mathbf{1^{\infty}}\}$ where $\mathbf{0^{\infty}},\mathbf{1^{\infty}}:L\rightarrow L$ are the constant maps identical to $0^{\infty}$ and $1^{\infty}$, respectively. Let $\Gamma$ be the semigroup generated by $\Delta$. The preceding proposition shows that $\Delta$ (and hence $\Gamma$) consists of continuous functions. In particular, the CCS $(L,\mathcal{P},\Gamma)$ satisfies the Extendibility Axiom. In contrast with Newton’s method, the algebraic structure of $\Delta$ is quite simple: composing two classifiers results in something similar to a Boolean logic gate. The topological structure is far more interesting. Intuitively, the crucial difference between Newton’s method and threshold classifiers is that the complexity of the former comes from depth: the semigroup is generated by a single map but its iterates are highly complex. The complexity of threshold classification comes from width: the semigroup has infinitely many generators, but their compositions are simple.
Intuitively, the closure of $\Delta$ consists of the set of all possible threshold classifiers on the real line, but there are two sorts: the ones that classify strict inequalities and those that classify $\leq$. The members of $\Delta$ are finite-precision approximations of classifiers that check all bits of information. But here it gets interesting: what is the difference, in terms of floating-point arithmetic, between $x<0.5$ and $x\leq 0.5$?
Suppose that $f_{a}^{+}$ represents the $\leq$ classifier for a target $a\in L$. We identify the scalar truth values with constant sequences, formally $f_{a}^{+}:L\to\{0^{\infty},1^{\infty}\}$ is defined by $f_{a}^{+}(x)=1^{\infty}$ if $x\leq_{\mathrm{lex}}a$ and $f_{a}^{+}(x)=0^{\infty}$ otherwise. Note that if $a$ is the constant $1^{\infty}$, then $f_{a}^{+}=\mathbf{1^{\infty}}$. Similarly, let $f_{a}^{-}$ be the strict inequality << classifier, i.e., $f_{a}^{-}(x)=1^{\infty}$ if $x<_{\mathrm{lex}}a$ and $f_{a}^{-}(x)=0^{\infty}$ otherwise. Note that if $a$ is the constant zero, then $f_{a}^{-}=\mathbf{0^{\infty}}$.
Proposition 2.9.
$f_{a}^{+},f_{a}^{-}\in\overline{\Delta}$ for all $a\in 2^{\mathbb{N}}$.
Proof.
First, we show that $f_{a}^{+}\in\overline{\Delta}$. If $a=1^{\infty}$, then $f_{a}^{+}=\mathbf{1^{\infty}}\in\Delta$. If $a$ is not identically $1$, we argue that the pointwise limit of the threshold classifiers on $w_{n}:=a|_{n}^{\frown}1$ (that is, the sequence obtained from appending a $1$ to the first $n$ bits of $a$) is precisely $f_{a}^{+}$. Specifically, for every $x\in L$, we intend to prove that $\lim_{n\to\infty}\phi_{w_{n}}(x)=f_{a}^{+}(x)$. Assume that $x>_{\mathrm{lex}}a$. Let $m$ be the least index at which the two sequences differ. Then $a(m)=0<1=x(m)$, and for all $n\geq m$, $w_{n}$ agrees with $a$ up to $m$. Crucially, $w_{n}(m)=0<1=x(m)$, which implies that $w_{n}<_{\mathrm{lex}}x|_{n+1}$, and hence $\phi_{w_{n}}(x)=0^{\infty}=f_{a}^{+}(x)$ for large enough $n$. If $x\leq_{\rm lex}a$, then $x|_{n+1}\leq_{\rm lex}w_{n}$ for all $n\in\mathbb{N}$. Hence, $\phi_{w_{n}}(x)=1^{\infty}=f_{a}^{+}(x)$ for all $n\in\mathbb{N}$.
Now, we prove that $f_{a}^{-}\in\overline{\Delta}$. If $a$ is the constant zero, then $f_{a}^{-}=\mathbf{0^{\infty}}\in\Delta$. Suppose that $a$ is not constantly zero; then we have two cases.
  1. (1)
    If $a$ is eventually zero ($a$ is often called a dyadic rational), that is $a=u^{\frown}1^{\frown}0^{\infty}$ (here denotes concatenation) for some finite $u$. Let $w_{n}:=u^{\frown}0^{\frown}1^{n}<_{\mathrm{lex}}a$. We claim that $\lim_{n\to\infty}\phi_{w_{n}}(x)=f_{a}^{-}(x)$. Assume that $x<_{\rm lex}a$. Then, $x|_{|w_{n}|}\leq_{\rm lex}w_{n}$ for large enough $n$. Hence, $\phi_{w_{n}}(x)=1^{\infty}=f_{a}^{-}(x)$ for large enough $n$. Now assume that $x\geq_{\rm lex}a$. Then, $w_{n}<_{\rm lex}a|_{|w_{n}|}\leq_{\rm lex}x|_{|w_{n}|}$ so $\phi_{w_{n}}(x)=0^{\infty}=f_{a}^{-}(x)$ for all $n\in\mathbb{N}$.
  2. (2)
    If $a$ is not eventually zero, enumerate the indices of all positive bits in $a$, $\{n\in\mathbb{N}:a(n)=1\}$, strictly increasingly as $\{n_{k}:k\in\mathbb{N}\}$ (this is possible as the former set is infinite by assumption). Let $w_{k}:=(a|_{n_{k}-1})^{\frown}0$; that is, $w_{k}$ is the result of flipping the $k$-th positive bit in $a$. Once again, observe that $w_{k}<_{\mathrm{lex}}a$ for all $k$. The crucial step follows from the fact that for any $x<_{\mathrm{lex}}a$, there is a large enough $K$ such that $x<_{\mathrm{lex}}w_{k}$ for all $k\geq K$.
The preceding proposition shows that the topological structure of deep computations can be very complicated. Indeed, $\overline{P_{n}\circ\Delta}$ contains the Split Cantor space for all $n\in\mathbb{N}$. (see Examples 3.3(3)).
2.5. Finite precision prefix test
In this subsection we present another example of a CCS with a more complicated set of deep computations. Let $L=2^{\mathbb{N}}$ and $\mathcal{P}=\{P_{n}:n\in\mathbb{N}\}$ where $P_{n}(x)=x(n)$ are the projection maps so clearly $L\subseteq\mathbb{R}^{\mathcal{P}}$ and $\mathcal{L}=L$ (same computation states structure as subsection 2.4). For each $w\in 2^{<\mathbb{N}}$, let $\psi_{w}:L\rightarrow L$ be the transition given by:
$\psi_{w}(x)=\begin{cases}1^{\infty},&\text{if }x|_{|w|}=w;\\ 0^{\infty},&\text{otherwise}.\end{cases}$
In other words, $\psi_{w}$ determines whether the first $|w|$ bits of a binary sequence is exactly $w$. Let $\Delta=\{\psi_{w}:w\in 2^{<\mathbb{N}}\}$ and $\Gamma$ be the semigroup generated by $\Delta$. Since the sets $\{x\in 2^{\mathbb{N}}:x|_{|w|}=w\}$ are open and closed in $2^{\mathbb{N}}$, then the transitions $\psi_{w}$ are all continuous. In particular, $(L,\mathcal{P},\Gamma)$ satisfies the Extendibility Axiom.
Let us analyze the set of deep computations of $\Delta$. The idea of these finite precision prefix tests $\psi_{w}$ is that they are approximating the equality relation on infinite binary sequences. For a given $a\in 2^{\mathbb{N}}$, let $\delta_{a}:L\rightarrow\{0^{\infty},1^{\infty}\}$ be the indicator function at $a$, i.e., $\delta_{a}(x)=1^{\infty}$ if $x=a$ and $\delta_{a}(x)=0^{\infty}$ otherwise.
Proposition 2.10.
$\delta_{a}\in\overline{\Delta}$ for all $a\in 2^{\mathbb{N}}$.
Proof.
Fix $a\in 2^{\mathbb{N}}$. Let $w_{n}:=a|_{n}$ for each $n\in\mathbb{N}$. We claim that $\lim_{n\rightarrow\infty}\psi_{w_{n}}(x)=\delta_{a}(x)$ for all $x\in L$. If $x=a$, then $x|_{|w_{n}|}=w_{n}$ for all $n$ and so $\psi_{w_{n}}(x)=1^{\infty}=\delta_{a}(x)$ for all $n$. If $x\neq a$, then $x|_{|w_{n}|}\neq w_{n}$ for large enough $n$. Hence, $\psi_{w_{n}}(x)=0^{\infty}=\delta_{a}(x)$ for large enough $n$. ∎
These equality tests $\delta_{a}$ are not all the deep computations. The other deep computation we are missing is the constant map $\mathbf{0^{\infty}}$.
Proposition 2.11.
$\mathbf{0^{\infty}}\in\overline{\Delta}$.
Proof.
To show that $\mathbf{0^{\infty}}\in\overline{\Delta}$ consider, for each $n\in\mathbb{N}$, ${w_{n}=1^{n}}^{\frown}0$, i.e., the string consisting of $n$ consecutive $1$s followed by a 0. If $x=1^{\infty}$, then $x|_{|w_{n}|}\neq w_{n}$ for all $n\in\mathbb{N}$. Hence, $\psi_{w_{n}}(x)=0^{\infty}$ for all $n\in\mathbb{N}$. If $x\neq 1^{\infty}$, let $N$ be the smallest such that $x(N)=0$. Then, $x|_{|w_{n}|}\neq w_{n}$ for all $n>N$. Hence, $\psi_{w_{n}}(x)=0^{\infty}$ for large enough $n$. ∎
In fact, $\overline{\Delta}=\Delta\cup\{\delta_{a}:a\in 2^{\mathbb{N}}\}\cup\{\mathbf{0^{\infty}}\}$ and this space is known as the Extended Alexandroff compactification of $2^{\mathbb{N}}$ (see Example 3.3(2)). One key topological property about this space is that $\mathbf{0^{\infty}}$ is not a $G_{\delta}$ point, i.e., $\{\mathbf{0^{\infty}}\}$ is not a countable intersection of open sets. Moreover, $\mathbf{0^{\infty}}$ is the only non-$G_{\delta}$ point. It is well-known that in a Hausdorff, first countable space every point is $G_{\delta}$. This shows that our space of deep computations is not first countable. This space also contains a discrete subspace of size continuum, namely $\{\delta_{a}:a\in 2^{\mathbb{N}}\}$.